Binomial Expansion- to determine the P of a particular set (combination) of events occuring in any order.
Question: What is the P of a couple having 2 boys and 4 girls out of 6 children?

Answer by stanbon(75887)

What is the P of a couple having 2 boys and 4 girls out of 6 children?
----
Binomial Problem with n = 6 and p(boy) = 1/2 = p(girl)
---
P(2boys in 6 births) = 6C2*(1/2)^2*(1/2)^4
-----
= 15*1/2^6
----
= 15/64
============
Cheers,
Stan H.
================